OpenStudy (anonymous):

(LCD) of 25/42 and 11/18 no answer just explainnn.. i factored (dnt know if thts a real word) 42 2x21 2x7 so 42 is 2x3x7 THEN 18 2x9 3x3 18 is 2x3x3 then they said. Find the LCM of 42 and 18. "Since we count the duplicate pairs of factors (circled below) only once, the LCM of 42 and 18 is 2×3×3×7, or 126. So, the LCD of the fractions is 126." why did they double the "3" and not the "2"? what made it 2x3x3x7 i need clarity for everything story of my life...

OpenStudy (gorv):

42=2*3*7 18=2*3*3 now u see 2*3 is common in both so it will be taken only once as it is common in both and 3 and 7 are remained there so lcm=2*3*3*7

OpenStudy (gorv):

@Anjxo ..gotta????

OpenStudy (anonymous):

yes thank you so much @gorv !x)